At a root 3highway, on ramp A car of mass 1.50*10^3Kg is stopped at a stop sign, waiting for a break in traffic before merging with the cars on the highway. A pickup of mass 2.00*10^3Kg comes up from behind and hit the stopped car. Assuming the collision is elastic, how fast was the pick-up going before collision straight head before the highway at 20.0ms^-1 just after the collision


1
Expert's answer
2021-03-09T15:26:56-0500
"v_{1f}=\frac{m_1-m_2}{m_1+m_2}v_{1i}+\frac{2m_2}{m_1+m_2}v_{2i}"

"v_{2i}=0"

Thus,

"20=\frac{2-1.5}{2+1.5}v_{1i}\\v_{1i}=140\frac{m}{s}"